# Mitigating Spurious Correlation via Distributionally Robust Learning with Hierarchical Ambiguity Sets

### v1

@@ 1 Introduction
In many instances of subpopulation shifts, the group indicator is spuriously correlated with the target label or response variable. For example, in the widely studied Waterbirds dataset (Sagawa et al., 2020), the target label (e.g., “Waterbird” or “Landbird”) is spuriously correlated with the background environment (e.g., water or land). As a result, ERM-based models tend to associate “Waterbird” primarily with water backgrounds, leading to significant performance degradation on minority groups, such as waterbirds appearing against land backgrounds. These vulnerabilities extend beyond controlled benchmarks, posing substantial risks in domains such as healthcare (Zech et al., 2018; Badgeley et al., 2019), fairness (Hashimoto et al., 2018; Buolamwini & Gebru, 2018; Obermeyer et al., 2019), and autonomous driving (Zhang et al., 2017).
@@
Over the past few years, a growing body of research has focused on mitigating these spurious correlations to ensure more reliable and stable model performance. Among the proposed methodologies, group distributionally robust optimization (Group DRO) (Sagawa et al., 2020) has emerged as a foundational approach. By partitioning the data into predefined groups and optimizing for the worst group loss, Group DRO effectively minimizes the model’s reliance on spurious correlations tied to specific subsets of data. This framework has inspired a wide range of subsequent methods, such as JTT (Liu et al., 2021), SSA (Nam et al., 2022), PG-DRO (Ghosal & Li, 2023), DISC (Wu et al., 2023) and GIC (Han & Zou, 2024), which commonly employ a two-step strategy: first identifying latent groups and then utilizing established robust training approaches—frequently Group DRO itself—to enhance model robustness.
While these methods have significantly advanced the field, they primarily focus on minimizing the worst-group loss under the assumption that each group’s training distribution reliably represents its true underlying distribution. However, this assumption often fails in practice—especially for minority groups with limited samples—where within-group distributional shifts naturally arise as a consequence of underrepresentation in the training data (Ben-Tal & Nemirovski, 2002; Devroye et al., 2013; Duchi & Namkoong, 2021). This limitation highlights the need for a more flexible and robust approach that accounts for uncertainty not only across groups, but also within them.
@@ 2 Related work
In parallel, a growing body of work addresses scenarios where group annotations are unavailable or prohibitively expensive, motivating the need to infer or approximate group structure. Building on Kirichenko et al. (2023), LaBonte et al. (2024) extends last-layer retraining to settings with minimal or no group annotations. Some methods employ limited validation sets to discover latent groups and then apply Group DRO (Sagawa et al., 2020) for robust optimization (Nam et al., 2022; Ghosal & Li, 2023). Indeed, the model’s loss (or its alternatives) often helps identify underrepresented subpopulations; for example, Nam et al. (2020); Liu et al. (2021); Qiu et al. (2023) use high-loss samples to recognize minority groups. Other approaches draw on diverse cues: Ahmed et al. (2021); Creager et al. (2021) infer groups by maximizing violations of the invariant risk minimization principle (Arjovsky et al., 2019), while Asgari et al. (2022) employs masking to reduce reliance on spurious features. Sohoni et al. (2020); Seo et al. (2022) instead cluster feature embeddings to discover latent groups and identify pseudo-attributes for debiasing. Likewise, Zhang et al. (2022) proposes a two-stage contrastive learning framework by aligning samples with the same class but different spurious attributes, and Wu et al. (2023) constructs a “concept bank” of candidate spurious attributes for robust partitioning. Another line of work identifies spurious features by comparing the training set with a carefully selected reference dataset (Han & Zou, 2024) or removes examples that disproportionately degrade worst-group accuracy (Jain et al., 2024).
Beyond these established techniques, emerging efforts explore new failure modes within the spurious correlation framework. For instance, recent works have examined imperfect group partitions (Zhou et al., 2021), multiple spurious features (Paranjape et al., 2023; Kim et al., 2024), low spurious signal strength (Mulchandani & Kim, 2025), and spurious-aware out-of-distribution detection (Zohrabi et al., 2025). In this context, our work addresses a critical yet underexplored issue: minority-group data, by nature extremely limited in size, may not faithfully reflect their underlying distributions—a natural but overlooked phenomenon that must be explicitly accounted for.
@@ 3 Preliminaries
Note that in frequently used DRO frameworks, the ambiguity set $\mathcal{Q}$ is often defined as a small neighborhood with respect to a standard (pseudo-)metric, such as the Wasserstein distance (Mohajerin Esfahani & Kuhn, 2018; Gao et al., 2024) and $f$-divergence (Namkoong & Duchi, 2016; Mehta et al., 2024).
@@ 4.1 Hierarchical Ambiguity Sets
The ambiguity set $\mathcal{Q}$ has a two-level hierarchy. The first level is controlled by the mixing proportion $\beta$. It accounts for uncertainty in the proportion of each subpopulation or group. Such uncertainty can arise, for example, if certain minority groups appear more frequently in evaluation settings than in the training set, thereby increasing their probability of occurrence and potentially amplifying spurious correlations if not properly addressed (Sugiyama & Storkey, 2006). At the second level, the distributional shift in each group is considered to capture within-group variability.
@@
The choice of $d_{1}$ is not critical because most reasonable metrics are equivalent. In the remainder of this paper, we set $\rho=\infty$.
For $d_{2}$, we consider the infinite-order Wasserstein distance for computational convenience. Recall the definition of the Wasserstein distance of order $p\in[1,\infty)$:
@@
where $A^{\epsilon}$ denotes the $\epsilon$-enlargement of $A$; see Givens & Shortt (1984).
The cost function is defined in a latent semantic space, which is more effective than defining it in the space of raw data (Zeiler & Fergus, 2014; Krizhevsky et al., 2017). Specifically, we employ a deep neural network $f^{\theta}$ of depth $L$, defined as
@@ 4.2 Algorithm
Intuitively, Theorem 4.1 shows that the worst-case risk over our hierarchical ambiguity set can be conservatively over-approximated by an adversarial perturbation problem in the latent space, where the inner maximization is weighted by the worst-case group proportions $\beta$. We therefore minimize the surrogate objective (9) via a coordinate-wise procedure, as detailed next.
@@ 5.2 Baselines
We compare our method to several representative baselines: ERM, Group DRO (Sagawa et al., 2020), JTT (Liu et al., 2021), CnC (Zhang et al., 2022), SSA (Nam et al., 2022), LISA (Yao et al., 2022), DFR (Kirichenko et al., 2023), PDE (Deng et al., 2024), and GIC (Han & Zou, 2024). These methods range from direct robust learning (e.g., Group DRO) to two-step pipelines that first infer group membership and then apply robust training (e.g., SSA, GIC). Detailed descriptions are provided in Appendix E.
@@ 6 Conclusion
We begin with a lemma adapted from Staib & Jegelka (2017), with minor adjustments to match our framework. This lemma provides an equivalent form for the inner supremum problem of DRO with a $W_{\infty}$-neighborhood, which is closely related to the representation (6) of $W_{\infty}$.
@@
(Staib & Jegelka, 2017, Proposition 3.1) Let $\theta$ be fixed model parameters, and let $c(\cdot,\cdot)$ be a metric on the input space $\mathcal{X}$. For any distribution $P$ on $\mathcal{X}\times\mathcal{Y}$ and for any $\epsilon\geq 0$,
@@ Appendix E Baseline Details
  • GIC (Han & Zou, 2024): Uses a two-step pipeline where group membership is partially inferred, then a robust optimization (e.g., Group DRO) is applied. Similar to LISA, it can incorporate tailored Mixup strategies depending on the dataset’s characteristics.
@@ Appendix F Implementation Details
For experiments involving our newly constructed datasets, we reimplemented both our proposed method and the relevant baselines. When certain baselines lacked reported results for a given dataset, we used the performance from Zhang et al. (2022) and Han & Zou (2024) if available; otherwise, we performed our own reimplementations under consistent settings. In particular, for the original CMNIST dataset, we reimplemented experiments for DFR and PDE, since their original papers did not include CMNIST results. In all other cases, we referenced performance metrics from each baseline’s primary source. All experiments were conducted on an NVIDIA GeForce RTX 3090 GPU.
@@ G.1 Selection of $\epsilon$
As is common in DRO problems, selecting the size of the ambiguity set $\epsilon$ is challenging. Since the extent of distributional shifts is generally unknown, many robust learning methods commonly rely on heuristic or even arbitrary choices in practice. We propose a simple data-driven procedure that partitions the training data using a one-dimensional t-SNE (Van der Maaten & Hinton, 2008) ordering. This avoids reliance on shifted or test-time data while still creating validation splits that mimic intra-group distribution shifts. Our design is inspired by prior work (Duchi & Namkoong, 2021), which used training-data partitioning to simulate distributional shifts.

### v2

@@ 1 Introduction
In many instances of subpopulation shifts, the group indicator is spuriously correlated with the target label or response variable. For example, in the widely studied Waterbirds dataset (Sagawa et al., 2020), the target label (e.g., “Waterbird” or “Landbird”) is spuriously correlated with the background environment (e.g., water or land). As a result, ERM-based models tend to associate “Waterbird” primarily with water backgrounds, leading to significant performance degradation on minority groups, such as waterbirds appearing against land backgrounds. These vulnerabilities extend beyond controlled benchmarks, posing substantial risks in real-world, high-stakes domains such as healthcare (Zech et al., 2018; Badgeley et al., 2019), fairness (Hashimoto et al., 2018; Buolamwini and Gebru, 2018; Obermeyer et al., 2019), and autonomous driving (Zhang et al., 2017).
@@
Over the past few years, a growing body of research has focused on mitigating these spurious correlations to ensure more reliable and stable model performance. Among the proposed methodologies, group distributionally robust optimization (Group DRO) (Sagawa et al., 2020) has emerged as a foundational approach. By partitioning the data into predefined groups and optimizing for the worst group loss, Group DRO effectively minimizes the model’s reliance on spurious correlations tied to specific subsets of data. This framework has inspired a wide range of subsequent methods, such as JTT (Liu et al., 2021), SSA (Nam et al., 2022), PG-DRO (Ghosal and Li, 2023), DISC (Wu et al., 2023) and GIC (Han and Zou, 2024), which commonly employ a two-step strategy: first identifying latent groups and then utilizing established robust training approaches—frequently Group DRO itself—to enhance model robustness.
While these methods have significantly advanced the field, they primarily focus on minimizing the worst-group loss under the assumption that each group’s training distribution reliably represents its true underlying distribution. However, this assumption often fails in practice—especially for minority groups with limited samples—where within-group distributional shifts naturally arise as a consequence of underrepresentation in the training data (Ben-Tal and Nemirovski, 2002; Devroye et al., 2013; Duchi and Namkoong, 2021). This limitation highlights the need for a more flexible and robust approach that accounts for uncertainty not only across groups, but also within them.
@@ 2 Related work
In parallel, a growing body of work addresses scenarios where group annotations are unavailable or prohibitively expensive, motivating the need to infer or approximate group structure. Building on Kirichenko et al. (2023), LaBonte et al. (2024) extends last-layer retraining to settings with minimal or no group annotations. Some methods employ limited validation sets to discover latent groups and then apply Group DRO (Sagawa et al., 2020) for robust optimization (Nam et al., 2022; Ghosal and Li, 2023). Indeed, the model’s loss (or its alternatives) often helps identify underrepresented subpopulations; for example, Nam et al. (2020); Liu et al. (2021); Qiu et al. (2023) use high-loss samples to recognize minority groups. Other approaches draw on diverse cues: Ahmed et al. (2021); Creager et al. (2021) infer groups by maximizing violations of the invariant risk minimization principle (Arjovsky et al., 2019), while Asgari et al. (2022) employs masking to reduce reliance on spurious features. Sohoni et al. (2020); Seo et al. (2022) instead cluster feature embeddings to discover latent groups and identify pseudo-attributes for debiasing. Likewise, Zhang et al. (2022) proposes a two-stage contrastive learning framework by aligning samples with the same class but different spurious attributes, and Wu et al. (2023) constructs a “concept bank” of candidate spurious attributes for robust partitioning. Another line of work identifies spurious features by comparing the training set with a carefully selected reference dataset (Han and Zou, 2024) or removes examples that disproportionately degrade worst-group accuracy (Jain et al., 2024).
Beyond these established techniques, emerging efforts explore new failure modes within the spurious correlation framework. For instance, recent works have examined imperfect group partitions (Zhou et al., 2021), multiple spurious features (Paranjape et al., 2023; Kim et al., 2024), low spurious signal strength (Mulchandani and Kim, 2025), and spurious-aware out-of-distribution detection (Zohrabi et al., 2025). Relatedly, recent work has shown that the presence of spurious correlations can itself constitute a failure mode for domain adaptation methods (Kim et al., 2026). In this context, our work addresses a critical yet underexplored issue: minority-group data, by nature extremely limited in size, may not faithfully reflect their underlying distributions—a natural but overlooked phenomenon that must be explicitly accounted for.
@@ 3 Preliminaries
Note that in frequently used DRO frameworks, the ambiguity set $\mathcal{Q}$ is often defined as a small neighborhood with respect to a standard (pseudo-)metric, such as the Wasserstein distance (Mohajerin Esfahani and Kuhn, 2018; Gao et al., 2024) and $f$-divergence (Namkoong and Duchi, 2016; Mehta et al., 2024).
@@ 4.1 Hierarchical Ambiguity Sets
The ambiguity set $\mathcal{Q}$ has a two-level hierarchy. The first level is controlled by the mixing proportion $\beta$. It accounts for uncertainty in the proportion of each subpopulation or group. Such uncertainty can arise, for example, if certain minority groups appear more frequently in evaluation settings than in the training set, thereby increasing their probability of occurrence and potentially amplifying spurious correlations if not properly addressed (Sugiyama and Storkey, 2006). At the second level, the distributional shift in each group is considered to capture within-group variability.
@@
The choice of $d_{1}$ is not critical because most reasonable metrics on $\Delta_{m-1}$ lead to similar behavior in our formulation. The constraint $d_{1}(\beta,\alpha)\leq\rho$ controls how much the test-time group proportions $\beta$ are allowed to deviate from the training mixture $\alpha$. When prior knowledge suggests that the test-time group proportions will not depart substantially from those observed in training (for example, they are expected to remain within a small neighborhood of $\alpha$), one may choose a finite $\rho$ to restrict attention to such mixtures. In the spurious-correlation setting we study, however, it is standard to assume no reliable prior on test-time group proportions and to evaluate robustness via worst-group accuracy, which can be viewed as examining performance under the extreme mixture that concentrates all probability mass on the worst-performing group. For this reason, in the remainder of this paper we instantiate our framework with $\rho=\infty$.
For $d_{2}$, we use a Wasserstein distance. Other options such as $f$-divergences (e.g., KL divergence or total variation distance) are in principle also possible, but they induce ambiguity sets with a structural limitation in our setting. Ambiguity sets defined via an $f$-divergence require candidate distributions to be absolutely continuous with respect to the empirical distribution; since the empirical distribution has finite support, such sets can only reweight the observed training samples. In contrast, Wasserstein distances do not impose absolute continuity and therefore allow support shifts, enabling the ambiguity set to include meaningful variations that may not appear in the training data. This choice is thus better aligned with the intra-group distributional shifts we aim to capture.
Among Wasserstein distances, we consider the infinite-order Wasserstein distance for computational convenience. Recall the definition of the Wasserstein distance of order $p\in[1,\infty)$:
@@
where $A^{\epsilon}$ denotes the $\epsilon$-enlargement of $A$; see Givens and Shortt (1984).
The cost function is defined in a latent semantic space, which is more effective than defining it in the space of raw data (Zeiler and Fergus, 2014; Krizhevsky et al., 2017). Specifically, we employ a deep neural network $f^{\theta}$ of depth $L$, defined as
@@ 4.2 Algorithm
Intuitively, Theorem 4.1 shows that the worst-case risk over our hierarchical ambiguity set can be conservatively over-approximated by an adversarial perturbation problem in the latent space, where the inner maximization is weighted by the worst-case group proportions $\beta$.
Remark (Tightness of the relaxation).
The upper bound in (9) relies on relaxing the inner maximization from
$\mathbb{E}_{P_{g}}\left[\sup_{x:\|z(x)-z(X)\|\leq\epsilon_{g}}\mathcal{L}\big(f^{\theta}_{L}(z(x)),Y\big)\right]$ $\leq\mathbb{E}_{P_{g}}\left[\sup_{z^{\prime}:\|z^{\prime}-z(X)\|\leq\epsilon_{g}}\mathcal{L}\big(f^{\theta}_{L}(z^{\prime}),Y\big)\right]$
This relaxation can be strict in general when the feature map $z(\cdot)$ is not surjective onto the $\epsilon_{g}$-ball. However, we expect this effect to be negligible in practice for embedding distributions arising in modern deep networks. It is commonly assumed that the pushforward distribution of $X$ under $z(\cdot)$ possesses a positive Lebesgue density on, or in a neighborhood of, its effective support.
Although raw images may concentrate near a low-dimensional manifold in input space and thus typically do not admit a Lebesgue density, the situation is different after they pass through several nonlinear layers. Empirically and conceptually, internal representations produced by deep networks behave as “thickened” manifolds and are typically modeled as distributions with nonzero volume in the ambient feature space.
Formally proving the existence of such a density for arbitrary deep embeddings is challenging. Nevertheless, under this standard modeling assumption, the relaxation above becomes nearly tight. The overlap between the $\epsilon_{g}$-ball and the image of $z(\cdot)$ is substantial, making the resulting upper bound in (9) a close approximation of the original hierarchical objective.
We therefore minimize the surrogate objective (9) via a coordinate-wise procedure, as detailed next.
@@ 5.2 Baselines
We compare our method to several representative baselines: ERM, Group DRO (Sagawa et al., 2020), JTT (Liu et al., 2021), CnC (Zhang et al., 2022), SSA (Nam et al., 2022), LISA (Yao et al., 2022), DFR (Kirichenko et al., 2023), PDE (Deng et al., 2024), and GIC (Han and Zou, 2024). These methods range from direct robust learning (e.g., Group DRO) to two-step pipelines that first infer group membership and then apply robust training (e.g., SSA, GIC). Detailed descriptions are provided in Appendix E.
@@ 6 Conclusion
Acknowledgments
This work was supported by the National Research Foundation of Korea (NRF) grant funded by the Korea government (MSIT) (No. RS-2023-00240861), a Korea Institute for Advancement of Technology (KIAT) grant funded by the Korea Government (MOTIE) (RS-2024-00409092, 2024 HRD Program for Industrial Innovation), and Institute of Information & Communications Technology Planning & Evaluation(IITP)-Global Data-X Leader HRD program grant funded by the Korea government (MSIT) (IITP-2024-RS-2024-00441244).
@@ Acknowledgments
We begin with a lemma adapted from Staib and Jegelka (2017), with minor adjustments to match our framework. This lemma provides an equivalent form for the inner supremum problem of DRO with a $W_{\infty}$-neighborhood, which is closely related to the representation (6) of $W_{\infty}$.
@@
(Staib and Jegelka, 2017, Proposition 3.1) Let $\theta$ be fixed model parameters, and let $c(\cdot,\cdot)$ be a metric on the input space $\mathcal{X}$. For any distribution $P$ on $\mathcal{X}\times\mathcal{Y}$ and for any $\epsilon\geq 0$,
@@ Appendix B Convergence Analysis of Algorithm 1
The convergence guarantee in Proposition B.1 is derived under the assumption that the feature map $z(\cdot)$ is fixed with respect to $\theta$. Under this assumption, the resulting optimization problem becomes convex in $\theta$, which allows us to establish the stated $O(1/\sqrt{T})$ convergence rate. Accordingly, Proposition B.1 should be interpreted as characterizing the behavior of Algorithm 1 in the idealized regime where the representation is fixed, rather than as a guarantee for the full nonconvex joint optimization of all network parameters.
@@ Appendix E Baseline Details
  • GIC (Han and Zou, 2024): Uses a two-step pipeline where group membership is partially inferred, then a robust optimization (e.g., Group DRO) is applied. Similar to LISA, it can incorporate tailored Mixup strategies depending on the dataset’s characteristics.
@@ Appendix F Implementation Details
For experiments involving our newly constructed datasets, we reimplemented both our proposed method and the relevant baselines. When certain baselines lacked reported results for a given dataset, we used the performance from Zhang et al. (2022) and Han and Zou (2024) if available; otherwise, we performed our own reimplementations under consistent settings. In particular, for the original CMNIST dataset, we reimplemented experiments for DFR and PDE, since their original papers did not include CMNIST results. In all other cases, we referenced performance metrics from each baseline’s primary source. All experiments were conducted on an NVIDIA GeForce RTX 3090 GPU, and we make our code available at https://github.com/Sung-Ho-Jo/hierarchical-dro.
@@ G.1 Selection of $\epsilon$
As is common in DRO problems, selecting the size of the ambiguity set $\epsilon$ is challenging. Since the extent of distributional shifts is generally unknown, many robust learning methods commonly rely on heuristic or even arbitrary choices in practice. We propose a simple data-driven procedure that partitions the training data using a one-dimensional t-SNE (Van der Maaten and Hinton, 2008) ordering. This avoids reliance on shifted or test-time data while still creating validation splits that mimic intra-group distribution shifts. Our design is inspired by prior work (Duchi and Namkoong, 2021), which used training-data partitioning to simulate distributional shifts.
@@ Appendix H Grad-CAM Results and Analysis
Appendix I Comparison with Standard Wasserstein DRO
To contextualize the performance gains achieved by our hierarchical formulation, we compare it against standard Wasserstein DRO, using the same latent-space cost function and identical training hyperparameters. The comparison is carried out on the Waterbirds and CelebA benchmarks under the minority group shift setting considered in our experiments, which reflects the primary distributional scenario targeted by our method. Figures LABEL:fig:wasserstein_waterbirds and LABEL:fig:wasserstein_celeba summarize the behavior of standard Wasserstein DRO across a wide range of Wasserstein radii.
Two structural limitations underlie the weak performance observed in these plots. First, standard Wasserstein DRO models uncertainty only through perturbations within a single Wasserstein ball around the empirical distribution, without incorporating any uncertainty over group proportions. In spurious-correlation settings, however, variation in group proportions constitutes a fundamental mode of distribution shift: a minority group that is scarcely observed during training may appear far more frequently at test time, and ignoring this possibility prevents standard Wasserstein DRO from addressing such mixture-level changes. Second, in spurious-correlation benchmarks, different groups have substantially different sample sizes. Our method is designed to reflect this by assigning a distinct within-group radius to each group ($\epsilon_{g}=\epsilon/\sqrt{n_{g}}$), so that smaller groups—whose empirical distributions are statistically less reliable—are assigned a larger perturbation radius. In contrast, standard Wasserstein DRO uses a single global radius and therefore cannot adapt to group-specific statistical uncertainty, which is crucial in the minority-shift setting.
For both Waterbirds and CelebA, we varied the Wasserstein radius over a broad range of values and report the corresponding worst-group accuracy curves below. Across all radii, standard Wasserstein DRO achieves significantly lower worst-group accuracy than our method, which attains $93.7\%$ and $72.1\%$ on Waterbirds and CelebA, respectively. These results highlight that perturbing the unconditional distribution alone is insufficient for capturing the types of shifts present in spurious-correlation settings.
(a) Performance of standard Wasserstein DRO across a wide range of radii.
These findings reinforce the main conclusion of our study: achieving robustness under the complex and realistic distribution shifts that arise in spurious-correlation benchmarks—particularly those involving minority-group shifts—requires modeling uncertainty over group proportions as well as group-specific conditional distributions.